# Compact LED-Based Displacement Sensing for Robot Fingers

### v2

@@ V-C Learning Algorithms
Force x (N) .0560 .0714 .990
Force y (N) .0533 .0687 .993
Force z (N) .0690 .0890 .560
Force magnitude (N) .0640 .0801 .978
@@ V-D Results and Discussion
The overall results of our sensor’s contact prediction are presented in Table II and Figure 8. In particular, Table II details the performance for each force and torque direction, as well as the overall force and torque magnitudes. Figure 8 visualizes the mean absolute error as a function of applied force and torque. In addition, Figure 7 provides a detailed performance analysis, showcasing all six force/torque dimensions across three exemplar contacts, each differing in the contact positions. We note again that the test data consisted of contacts never seen in training.
We observe that the sensor provides accurate measurements in both force and torque for the x and y axes, but less so for the z axis. Noise levels are generally low, and predictions are accurate in terms of overall force magnitude, but can occasionally suffer from accuracy at higher forces in individual DOFs. The model was evaluated primarily on the 0-5 N range of forces, but experimentation showed that training on data that extended beyond this range proved useful for the models learning.
We believe that one likely source of error is hysteresis introduced by the PDMS and quantified in Sec. IV. This is particularly prevalent with larger forces, where signals will take some time to settle back to their baselines, which introduces noise to forces done before the resettling. It also introduces noise into the zero-force data.
@@ VI Conclusion and Future Work
In this paper, we introduce a low-cost, compliant sensor designed to be easily integrated in a robot finger and provide information on the displacement induced by contact with the world. Our method relies on sensing the displacement between two rigid plates, each mounted with LED emitters and receivers. We analyzed such an LED-LED pair in isolation, and found it to be highly sensitive to relative displacements of the emitter and the receiver. We thus designed a sensor with six LEDs acting as emitters and 24 LEDs acting as receivers, distributed between the top and bottom plates of the sensor.
To characterize the information contained in the raw signals from our sensor, we tested wether it is possible to reconstruct overall finger wrench information, using a commercial force/torque sensor ars ground truth. Our results showed that a simple supervised learning model trained on data from the 24 receiver LEDs can learn to map changes in these signals due to displacement of the flexure to external forces and torques applied between the two plates. The sensor output was particularly accurate for force/torques along/around the x and y axes, with additional work needed to improve sensitivity for the z axis.
@@
Future work of the sensor falls under two categories: further improvement of the sensor, and implementation within actual use robot hands. Sensor improvement in turn will have mechanical, electronic, and learning components. Mechanically, we would like to reduce hysteresis, simplify the manufacturing, and improve the consistency between sensors. We would also like to reduce the stiffness in the z-axis so that we get more sensitivity in that direction. Electronically, we will continue to look for sources of noise in LED signals. We also want to minimize the form factor as much as possible, so further experiments would include use of fewer LEDs in each cluster (or potentially even fewer clusters). Learning wise, we will further tune the model as we gather more data for Z axis and explore alternate architectures with both time series and single-snapshot models.
@@
We look to take advantage of the features of the sensor which make it particularly attractive for integration in robot hands. In particular, the form factor and ease of integration (the sensor directly connects to a micro-controller board with no additional amplification electronics) are highly appealing. Figure 9 shows a possible design of an anthropomorphic hand with one of our sensors at the base of the distal phalanx of each finger. Such integration will also allow us to test the hypothesis that slight compliance in the sensing element can prove beneficial, or at least non-detrimental, to dexterous manipulation.
As motor learning for dexterous manipulation makes continuous advances, we believe that sensing will prove key for future performance of robot hands. Having a multitude of sensors at our disposal, each with various strengths, will increase the ability of combining the right sensor with the right motor learning method to achieve novel capabilities. We believe that the sensor presented in this paper represents a step in this direction.

### v3

@@ V-C Learning Algorithms
Force x (N) 0.06 0.07 .990
Force y (N) 0.05 0.07 .993
Force z (N) 0.07 0.09 .560
Force magnitude (N) 0.06 0.08 .978
@@ V-D Results and Discussion
The overall results of our sensor’s contact prediction are presented in Table II and Figure 8. In particular, Table II details the performance for each force and torque direction, as well as the overall force and torque magnitudes. Figure 8 visualizes the mean absolute error as a function of applied force and torque. In addition, Figure 7 provides a detailed performance analysis, showcasing all six force/torque dimensions across three exemplar contacts, each differing in the contact positions. We note again that the test data consisted of contacts never seen in training. Moreover, as we are leveraging LEDs as receivers, our model can operate with inputs streaming in at 250 Hz.
We observe that the sensor provides accurate measurements in both force and torque for the x and y axes, but less so for the z axis. Predictions are accurate in terms of overall force magnitude, but can occasionally suffer from inaccuracy at higher forces in individual DOFs. The model was evaluated primarily on the 0-5 N range of forces, but experimentation showed that training on data that extended beyond this range proved useful for the models learning.
We believe that one likely source of error is hysteresis introduced by the PDMS and quantified in Sec. IV. This is particularly prevalent with larger forces, where signals will take some time to settle back to their baselines, which introduces noise to forces done before the resettling, and also introduces noise into the zero-force data.
@@ VI Conclusion and Future Work
In this paper, we introduce a low-cost, compliant sensor designed to be easily integrated in a robot finger and provide information on the displacement induced by contact with the world. Our method relies on sensing the displacement between two rigid plates, each mounted with LED emitters and receivers. We analyzed such an LED-LED pair in isolation, and found it to be highly sensitive to relative displacements of the emitter and the receiver. We thus designed a sensor with six LEDs acting as emitters and 24 LEDs acting as receivers, distributed between the top and bottom plates.
To characterize the information contained in the raw signals from our sensor, we tested whether it is possible to reconstruct overall finger wrench information, using a commercial force/torque sensor as ground truth. Our results showed that a simple supervised learning model trained on data from the 24 receiver LEDs can learn to map changes in these signals due to displacement of the flexure to external forces and torques applied between the two plates. The sensor output was particularly accurate for force/torques along/around the x and y axes, with additional work needed to improve sensitivity for the z axis.
@@
Future work of the sensor falls under two categories: further improvement of the sensor, and integration within functional robot hands. Sensor improvement in turn will have mechanical, electronic, and learning components. The focus along all these domains will be to make a more robust, consistent, and easy to manufacture device. Although we would like to eliminate hysteresis, ultimately some level of time dependent non-linearity will have to be accepted. Therefore, in making decisions different flexure designs in the future, we will prioritize robustness, ease of manufacture, and consistency. Electronically, we will continue to look for sources of noise in LED signals. We also want to minimize the form factor as much as possible, so further experiments would include use of fewer LEDs in each cluster (or potentially even fewer clusters). Learning wise, we will further tune the model as we gather more data for z axis and explore alternate architectures with both time series and single-snapshot models.
@@
We look to take advantage of the features of the sensor which make it particularly attractive for integration in robot hands. In particular, the form factor and ease of integration (the sensor directly connects to a micro-controller board with no additional amplification electronics) are highly appealing. Figure 9 shows a possible design of an anthropomorphic hand with one of our sensors at the base of the distal phalanx of each finger. Such integration will also allow us to test the hypothesis that slight compliance in the sensing element can prove beneficial to dexterous manipulation.
As motor learning for dexterous manipulation makes continuous advances, we believe that sensing will prove key for future performance of robot hands. Having a multitude of sensors at our disposal, each with various strengths, will increase the ability of combining the right sensor with the right motor learning method to achieve novel capabilities. In particular, sensors that provide information on overall contact forces should be ubiquitous, commodity technology, available in any lab by the dozen, ready to be quickly integrated in any hand design, as opposed to expensive, fragile pieces connecting to complex external electronics and thus used rarely, if at all. We believe that the design presented in this paper represents a step in this direction.